# Composed Multi-modal Retrieval: A Survey of Approaches and Applications

### v1

@@ 6 Applications
At the application level, we categorize and summarize existing multimodal retrieval applications based on specific scenarios, including fashion images, natural images, videos, remote sensing images, pedestrian re-identification, and skeletal diagrams. The specific contexts and challenges faced by each task in these different applications are thoroughly summarized, with the hope of inspiring exploration in other yet-to-be-explored scenarios.
@@ 6.1 Composed Fashion Image Retrieval
Traditional fashion image retrieval primarily relies on simple image search or keyword-based search. However, these methods often fail to meet user needs when searching for specific fashion items with complex attributes such as color, style, and material. The main goal of Composed Fashion Image Retrieval (CFIR) is to achieve more accurate and personalized fashion searches by combining both images and texts. Users can input a reference image (e.g., a photo of a clothing item) along with a textual modification (e.g., ”long-sleeve version”), allowing the system to understand both the visual and semantic information, leading to more refined retrieval.
CFIR can be widely applied on e-commerce platforms, helping users quickly find products that meet specific needs, thus improving the convenience and accuracy of the purchase. Users can not only search by image but also refine their queries with textual descriptions, improving the efficiency and quality of search results. This increases the precision of online shopping, making it easier for users to find products that match their requirements, boosting customer satisfaction, and increasing overall user engagement. In addition, for retailers and brands, CFIR enables more accurate product recommendations based on user preferences while also supporting fashion trend analysis and inventory management.
@@ 6.1.1 Benchmark Datasets
Shoes [84] is collected from the Internet, specifically from like.com, a shopping website that aggregates product data from a wide range of e-Commerce sources. They crawl 10,751 pairs of shoe images with relative expressions that describe fine-grained visual differences. Typically, 10,000 samples are used for training and 4,658 samples are used for evaluation.
CSS [18] consists of 32K queries (16K for training and 16K for test), such as ’make yellow sphere small’ that serve as modification text for the images synthesized in a 3-by-3 grid scene. Although it is a relatively simple dataset, CSS has the benefit of facilitating carefully controlled experiments.
CFQ [17] is collected from a set of images from online advertisements from various fashion brands and retailers that feature one dress, usually worn by a model. It has explicit positive and negative labels for all images in the same category as responses to each query. The data set contains 334 images with 3 categories of dress, dress, and sundress and 40 queries.
@@ 6.1.2 Results and Analysis
TABLE I: Comparison of Different Methods on FashionIQ Dataset(Original Split)
@@
JSVM20’ECCV MobileNet LSTM $10.70$ $25.90$ $12.00$ $27.10$ $13.00$ $26.90$ $11.90$ $26.63$
TRACE21’AAAI ResNet50 GRU $22.70$ $44.91$ $20.80$ $40.80$ $24.22$ $49.80$ $22.57$ $46.19$
ComposeAE21’WACV ResNet18 BERT $10.77$ $28.29$ $9.96$ $25.14$ $12.74$ $30.79$ $11.80$ $29.40$
CoSMo21’CVPR ResNet50 LSTM $21.39$ $44.45$ $16.90$ $37.49$ $21.32$ $46.02$ $19.87$ $42.62$
FashionVLP22’CVPR ResNet50 BERT $26.77$ $53.2$ $22.67$ $46.22$ $28.51$ $57.47$ $25.98$ $52.30$
Combiner22’CVPR-D CLIP(RN50) CLIP $31.63$ $56.67$ $36.36$ $51.86$ $38.19$ $62.42$ $35.39$ $59.03$
MultiColSAP22’TMM ResNet50 LSTM $30.74$ $55.92$ $26.05$ $50.64$ $34.42$ $61.14$ $30.4$ $55.9$
CLIP4Cir122’CVPR-W CLIP(RN50) CLIP $33.81$ $59.40$ $39.99$ $60.45$ $41.41$ $65.37$ $38.32$ $61.74$
Progressive22’SIGIR ResNet50 BERT $29.00$ $53.94$ $35.43$ $58.88$ $39.16$ $64.56$ $34.53$ $59.13$
Progressive22’SIGIR CLIP(ViT-B/32) BERT $33.60$ $58.90$ $39.45$ $61.78$ $43.96$ $68.33$ $39.02$ $63.00$
ARTEMIS22’ICLR ResNet50 BiGRU $25.68$ $51.05$ $21.57$ $44.13$ $28.59$ $55.06$ $25.28$ $50.08$
ComqueryFormer23’TMM Swin-T BERT $28.85$ $55.38$ $25.64$ $50.22$ $33.61$ $60.48$ $29.37$ $55.36$
TransAgg(LaionCIR)23’BMVC BLIP BLIP $27.67$ $49.38$ $32.83$ $52.31$ $35.70$ $58.08$ $32.07$ $53.26$
CLIP-CD23’AI CLIP(RN50) CLIP $37.68$ $62.62$ $42.44$ $63.74$ $45.33$ $67.72$ $41.82$ $64.79$
CLIP4Cir223’TOMM CLIP(RN50) CLIP $37.67$ $63.16$ $39.87$ $60.84$ $44.88$ $68.59$ $40.80$ $64.20$
BLIP4CIR24’WACV BLIP BLIP $40.65$ $66.34$ $37.49$ $60.06$ $43.60$ $67.77$ $39.34$ $63.85$
BLIP4CIR+Bi24’WACV BLIP BLIP $42.09$ $67.33$ $41.76$ $64.28$ $46.61$ $70.32$ $43.49$ $67.31$
SPIRIT24’TOMM CLIP(RN50) CLIP $39.86$ $64.30$ $47.67$ $71.10$ $44.11$ $65.60$ $43.88$ $67.20$
DQU-CIR24’SIGIR CLIP(ViT-H/14) $51.90$ $74.37$ $53.57$ $73.21$ $58.48$ $79.23$ $54.65$ $75.60$
FashionViL24’ECCV ResNet50 BERT $33.47$ $59.94$ $25.17$ $50.39$ $34.98$ $60.79$ $31.21$ $57.04$
BLIP4CIR224’TMLR BLIP(ViT-B) BLIP $43.78$ $71.34$ $50.15$ $71.25$ $55.23$ $76.80$ $51.17$ $73.13$
CompoDiff23’TMLR CLIP(ViT-L/14) $35.53$ $49.56$ $40.88$ $53.06$ $41.15$ $54.12$ $39.05$ $52.34$
CompoDiff23’TMLR CLIP(ViT-G/14) $38.39$ $51.03$ $41.68$ $56.02$ $45.70$ $57.32$ $39.81$ $51.90$
SDQUR24’TCSVT BLIP2 $49.93$ $73.33$ $56.87$ $76.50$ $59.66$ $79.25$ $55.49$ $76.36$
AlRet24’TMM CLIP LSTM $35.75$ $60.56$ $37.02$ $60.05$ $42.25$ $67.52$ $38.20$ $62.82$
NSFSE24’TMM ResNet50 BiGRU $29.62$ $54.41$ $22.96$ $45.93$ $31.08$ $57.01$ $27.84$ $52.39$
NSFSE24’TMM ResNet152 BiGRU $31.12$ $55.73$ $24.58$ $45.85$ $31.93$ $58.37$ $29.17$ $53.24$
MANME24’TCSVT ResNet50 BiGRU $31.26$ $57.66$ $26.37$ $47.94$ $36.33$ $59.31$ $29.95$ $54.90$
@@
PALAVAR22’ECCV CLIP(ViT-B/32) $17.25$ $35.94$ $21.49$ $37.05$ $20.55$ $38.76$ $19.76$ $37.25$
Pic2Word23’CVPR CLIP(ViT-L/14) $20.00$ $40.20$ $26.20$ $43.60$ $27.90$ $47.40$ $24.70$ $43.70$
SEARLE23’ICCV CLIP(ViT-B/16) $18.54$ $38.51$ $24.44$ $41.61$ $25.70$ $46.46$ $22.89$ $42.53$
SEARLE-XL-OTI23’ICCV CLIP(ViT-L/14) $21.57$ $44.47$ $30.37$ $47.49$ $30.9$ $51.76$ $27.61$ $47.9$
SEARLE-XL23’ICCV CLIP(ViT-L/14) $20.48$ $43.13$ $26.86$ $45.58$ $29.32$ $49.97$ $25.56$ $46.23$
MTCIR23’Arxiv CLIP(ViT-L/14) $28.11$ $51.12$ $38.63$ $58.51$ $39.42$ $62.68$ $35.39$ $57.44$
VQA4CIR(+CLIP4CIR)23’Arxiv LLaMA + LLaVA $40.91$ $65.13$ $45.62$ $65.68$ $49.21$ $71.22$ $45.24$ $67.34$
VQA4CIR(+SPRC)23’Arxiv LLaMA + LLaVA $49.18$ $73.06$ $56.79$ $74.52$ $59.67$ $79.30$ $55.21$ $75.62$
KEDs24’CVPR CLIP(ViT-L/14) $21.7$ $43.8$ $28.9$ $48.0$ $29.9$ $51.9$ $26.8$ $47.9$
PM24’Arxiv CLIP(ViT-L/14) $27.1$ $43.8$ $21.4$ $41.7$ $28.9$ $47.3$ $25.8$ $44.2$
FTI4CIR24’SIGIR CLIP(ViT-L/14) $24.39$ $47.84$ $31.35$ $50.59$ $32.43$ $54.21$ $29.39$ $50.88$
CIReVL24’ICLR CLIP(ViT-B/32) $25.29$ $46.36$ $28.36$ $47.84$ $31.21$ $53.85$ $28.29$ $49.35$
CIReVL24’ICLR CLIP(ViT-L/14) $24.79$ $44.76$ $29.49$ $47.40$ $31.36$ $53.65$ $28.55$ $48.57$
CIReVL24’ICLR CLIP(ViT-G/14) $27.07$ $49.53$ $33.71$ $51.42$ $35.80$ $56.14$ $32.19$ $52.36$
LinCIR24’CVPR CLIP(ViT-L/14) $20.92$ $42.44$ $29.10$ $46.81$ $28.81$ $50.18$ $26.28$ $46.49$
LinCIR24’CVPR CLIP(ViT-H/14) $29.80$ $52.11$ $36.90$ $57.75$ $42.07$ $62.52$ $36.26$ $57.46$
LinCIR24’CVPR CLIP(ViT-G/14) $38.08$ $60.88$ $46.76$ $65.11$ $50.48$ $71.09$ $45.11$ $65.69$
iSEARLE-XL-OTI24’Arxiv CLIP(ViT-L/14) $24.19$ $45.12$ $31.8$ $50.2$ $31.72$ $53.29$ $29.24$ $49.54$
iSEARLE-XL24’Arxiv CLIP(ViT-L/14) $22.51$ $46.36$ $28.75$ $47.84$ $31.31$ $52.68$ $27.52$ $48.96$
RTD(SEARLE)24’Arxiv CLIP(ViT-B/32) $20.72$ $43.13$ $26.69$ $44.31$ $26.67$ $48.75$ $24.7$ $45.4$
RTD((LinCIR)24’Arxiv CLIP(ViT-B/32) $19.98$ $41.75$ $23.65$ $42.74$ $24.73$ $46.56$ $22.79$ $43.68$
RTD(SEARLE)24’Arxiv CLIP(ViT-B/32) $23.2$ $47.25$ $32.63$ $50.39$ $32.18$ $54.56$ $29.34$ $50.73$
RTD((LinCIR)24’Arxiv CLIP(ViT-L/14) $24.49$ $48.24$ $32.83$ $50.44$ $33.4$ $54.56$ $30.24$ $51.08$
ISA(Sym)24’ICLR BLIP $24.69$ $43.88$ $30.79$ $50.05$ $33.91$ $53.65$ $29.79$ $49.19$
ISA(Asy)24’ICLR EfficientNet+BLIP $25.33$ $46.26$ $30.03$ $48.58$ $33.45$ $53.8$ $29.6$ $49.54$
LDRE24’SIGIR CLIP(ViT-B/32) $19.97$ $41.84$ $27.38$ $46.27$ $27.07$ $48.78$ $24.81$ $45.63$
LDRE24’SIGIR CLIP(ViT-L/14) $22.93$ $46.76$ $31.04$ $51.22$ $31.57$ $53.64$ $28.51$ $50.54$
LDRE24’SIGIR CLIP(ViT-G/14) $26.11$ $51.12$ $35.94$ $58.58$ $35.42$ $56.67$ $32.49$ $55.46$
Context-I2W24’AAAI CLIP(ViT-L/14) $23.1$ $45.3$ $29.7$ $48.6$ $30.6$ $52.9$ $27.8$ $48.9$
Slerp24’ECCV CLIP(ViT-B/32) $20.53$ $41.00$ $23.75$ $40.92$ $26.98$ $46.77$ $23.75$ $42.90$
Slerp+TAT24’ECCV CLIP(ViT-B/32) $19.24$ $42.14$ $23.06$ $41.95$ $26.57$ $47.78$ $22.96$ $43.96$
Slerp+TAT24’ECCV CLIP(ViT-L/14) $23.35$ $45.12$ $29.94$ $46.47$ $31.97$ $51.20$ $28.32$ $47.00$
Slerp+TAT24’ECCV BLIP(ViT-L/16) $29.15$ $50.62$ $32.14$ $51.62$ $37.02$ $57.73$ $32.77$ $53.32$
HyCIR(CC3M+syn)24’Arxiv CLIP $19.98$ $40.8$ $27.62$ $44.94$ $28.14$ $47.67$ $25.25$ $44.47$
HyCIR(CC3M+syn)24’Arxiv BLIP $18.88$ $34.5$ $22.52$ $37.58$ $22.13$ $40.33$ $21.18$ $37.47$
@@
CompoDiff(ST18M)23’TMLR CLIP(ViT-L/14) $32.24$ $46.27$ $37.69$ $49.08$ $38.12$ $50.57$ $36.02$ $48.64$
CompoDiff(ST18M)23’TMLR CLIP(ViT-G/14) $37.78$ $49.10$ $41.31$ $55.17$ $44.26$ $56.41$ $39.02$ $51.71$
PTG(+SPRC)24’Arxiv $31.10$ $51.90$
CASE24’AAAI BLIP $47.44$ $69.36$ $48.48$ $70.23$ $50.18$ $72.24$ $48.79$ $70.68$
VDG(+DF’se)24’CVPR BLIP $47.10$ $69.10$ $49.95$ $69.96$ $53.90$ $74.35$ $50.32$ $71.14$
VDG(+FIQ’se)24’CVPR BLIP $47.89$ $69.81$ $51.36$ $71.08$ $53.29$ $74.65$ $50.85$ $71.85$

### v2

@@ 6 Applications
At the application level, as shown in Fig. 7, we categorize existing composed multimodal retrieval applications based on specific scenarios, including fashion images, natural images, videos, remote sensing images, person re-identification, skeletal images, and interactive conversation. The specific contexts and challenges faced by each task in these different applications are thoroughly summarized, with the hope of inspiring exploration in other yet-to-be-explored scenarios.
To assess the effectiveness of composed multimodal retrieval models across various applications, several standard evaluation metrics are commonly employed. Among them, Recall@K is the most widely used. It measures the proportion of relevant items retrieved among the top-K results. Due to the presence of potential false negatives in annotations, high values of K (e.g., 10, 50) are often used to mitigate incomplete labeling effects. We report average Recall@1, Recall@10, and Recall@50 following standard practice. Another commonly used metric is Mean Average Precision at K (mAP@K), which accounts for both precision and the rank of relevant items within the top-K results. It is particularly helpful when evaluating models that must not only retrieve relevant samples but also rank them effectively. These metrics provide a comprehensive view of both retrieval accuracy and ranking quality, and they are used selectively across different datasets and tasks according to their characteristics.
Figure 7: Existing composed multi-modal retrieval (CMR) applications can be categorized based on specific scenarios, including fashion images, natural images, videos, remote sensing images, person re-identification, skeletal images, and interactive conversation..
@@ 6.1 Composed Fashion Image Retrieval
Traditional fashion image retrieval primarily relies on simple image search or keyword-based search. However, these methods often fail to meet user needs when searching for specific fashion items with complex attributes such as color, style, and material. The main goal of Composed Fashion Image Retrieval (CFIR) is to achieve more accurate and personalized fashion searches by combining both images and texts. CFIR has broad applications in e-commerce. By combining image-based search with textual refinement, it improves search efficiency and helps users locate products that better match their preferences. For example, users can input a reference image (e.g., a photo of a clothing item) along with a textual modification (e.g., ”long-sleeve version”). This leads to higher customer satisfaction and greater user engagement. For retailers and brands, CFIR supports personalized product recommendations, fashion trend analysis, and more effective inventory management.
@@ 6.1.1 Benchmark Datasets
Shoes [28] is collected from the Internet, specifically from like.com, a shopping website that aggregates product data from a wide range of e-commerce sources. They crawl 10,751 pairs of shoe images with relative expressions that describe fine-grained visual differences. Typically, 10,000 samples are used for training, and 4,658 samples are used for evaluation.
There are also some datasets that are not commonly used, including CFQ [32], UT-Zap50k [102, 103], FashionGen [104], FACAD [105], and PolyvoreOutfits [106].
@@ 6.1.2 Results and Analysis
We evaluate the performance of various methods on three widely used benchmarks for composed fashion image retrieval: FashionIQ, Fashion200K, and Shoes, covering diverse types of textual modifications and attribute granularities. For the FashionIQ dataset, we follow two commonly adopted evaluation protocols: Original Split and VAL Split, as shown in Table I and Table II, respectively. The VAL Split, introduced in early CMR studies, constructs a reduced candidate pool within the validation set, resulting in a narrower search space and relatively simpler task. In contrast, the Original Split leverages the full candidate pool provided by the dataset, substantially increasing the retrieval difficulty and typically leading to lower Recall@K scores. In recent years, the Original Split has been increasingly adopted as a more rigorous benchmark for evaluating model generalization under large-scale open-domain settings.
Under the Original Split, early supervised methods typically employed ResNet + RNN architectures. For instance, JSVM only achieved an average R@50 of 26.6%. With the emergence of large-scale vision-language pretraining, models such as CLIP and BLIP have been progressively introduced into compositional retrieval tasks. CLIP4CIR2 and BLIP4CIR2 achieved average R@50 scores of 64.2% and 73.1%, respectively, demonstrating significant performance improvements. More recent methods, such as DQU-CIR and SDQUR, adopt non-explicit fusion architectures. Specifically, DQU-CIR performs modality fusion at the data level, while SDQUR leverages BLIP2’s Q-Former for diverse query generation and uncertainty modeling. These models achieve average R@50 scores of 75.6% and 76.4%, respectively. These results highlight the importance of backbone strength and deep fusion mechanisms in advancing compositional retrieval performance.
In the zero-shot learning (ZSL-CMR) setting, models do not involve task-specific training and rely entirely on the inherent cross-modal alignment capabilities of VLMs. For example, PrediCIR, using a ViT-G/14 backbone, achieves an average R@10 of 47.2%, comparable to or even outperforming several early supervised models. Other approaches such as FTI4CIR (pseudo-token based), LinCIR (self-supervised in the text domain), Slerp (spherical interpolation), and CIReVL (LLM-based retrieval) explore different composition paradigms, contributing to the growing diversity and maturity of the ZSL-CMR paradigm.
Under the semi-supervised learning (SSL-CMR) setting, methods such as VDG and CASE generate large-scale pseudo-triplets by using VLMs and LLMs to produce fine-grained textual differences for visually similar image pairs. Built on BLIP backbones, VDG and CASE achieve average R@10 scores of 50.85% and 48.79%, respectively, outperforming all existing ZSL methods and many supervised baselines. These results suggest that high-quality, automatically constructed pseudo-supervision can provide rich and effective learning signals. Furthermore, methods like CompoDiff extend this paradigm by directly generating target images via diffusion models, enabling more complex compositional scene synthesis.
For the Fashion200K and Shoes datasets (Table III), despite differing task emphases, we observe similar trends to FashionIQ. Fashion200K features broader category diversity; the R@50 improved from 63.8% (TIRG) to 87.8% (DQU-CIR). In the Shoes dataset, which requires finer attribute matching, performance increased from 75.8% (VAL) to 88.3% (SDQUR). These results indicate that stronger vision backbones (e.g., ViT, BLIP2), coupled with more expressive multimodal fusion modules (e.g., Q-Former, set-level alignment), consistently lead to improved retrieval accuracy.
TABLE I: Comparison of Different Methods on FashionIQ Dataset (Original Split).
@@
JSVM20’ECCV [69] MobileNet LSTM $10.70$ $25.90$ $12.00$ $27.10$ $13.00$ $26.90$ $11.90$ $26.63$
TRACE21’AAAI [50] ResNet50 GRU $22.70$ $44.91$ $20.80$ $40.80$ $24.22$ $49.80$ $22.57$ $46.19$
ComposeAE21’WACV [38] ResNet18 BERT $10.77$ $28.29$ $9.96$ $25.14$ $12.74$ $30.79$ $11.80$ $29.40$
CoSMo21’CVPR [37] ResNet50 LSTM $21.39$ $44.45$ $16.90$ $37.49$ $21.32$ $46.02$ $19.87$ $42.62$
FashionVLP22’CVPR [56] ResNet50 BERT $26.77$ $53.2$ $22.67$ $46.22$ $28.51$ $57.47$ $25.98$ $52.30$
Combiner22’CVPR-D [33] CLIP(RN50) CLIP $31.63$ $56.67$ $36.36$ $51.86$ $38.19$ $62.42$ $35.39$ $59.03$
MultiColSAP22’TMM [124] ResNet50 LSTM $30.74$ $55.92$ $26.05$ $50.64$ $34.42$ $61.14$ $30.4$ $55.9$
CLIP4Cir122’CVPR-W [34] CLIP(RN50) CLIP $33.81$ $59.40$ $39.99$ $60.45$ $41.41$ $65.37$ $38.32$ $61.74$
Progressive22’SIGIR [237] ResNet50 BERT $29.00$ $53.94$ $35.43$ $58.88$ $39.16$ $64.56$ $34.53$ $59.13$
Progressive22’SIGIR [237] CLIP(ViT-B/32) BERT $33.60$ $58.90$ $39.45$ $61.78$ $43.96$ $68.33$ $39.02$ $63.00$
ARTEMIS22’ICLR [129] ResNet50 BiGRU $25.68$ $51.05$ $21.57$ $44.13$ $28.59$ $55.06$ $25.28$ $50.08$
ComqueryFormer23’TMM [238] Swin-T BERT $28.85$ $55.38$ $25.64$ $50.22$ $33.61$ $60.48$ $29.37$ $55.36$
TransAgg(LaionCIR)23’BMVC [91] BLIP BLIP $27.67$ $49.38$ $32.83$ $52.31$ $35.70$ $58.08$ $32.07$ $53.26$
CLIP-CD23’AI [36] CLIP(RN50) CLIP $37.68$ $62.62$ $42.44$ $63.74$ $45.33$ $67.72$ $41.82$ $64.79$
CLIP4Cir223’TOMM [35] CLIP(RN50) CLIP $37.67$ $63.16$ $39.87$ $60.84$ $44.88$ $68.59$ $40.80$ $64.20$
BLIP4CIR24’WACV [93] BLIP BLIP $40.65$ $66.34$ $37.49$ $60.06$ $43.60$ $67.77$ $39.34$ $63.85$
BLIP4CIR+Bi24’WACV [93] BLIP BLIP $42.09$ $67.33$ $41.76$ $64.28$ $46.61$ $70.32$ $43.49$ $67.31$
SPIRIT24’TOMM [61] CLIP(RN50) CLIP $39.86$ $64.30$ $47.67$ $71.10$ $44.11$ $65.60$ $43.88$ $67.20$
DQU-CIR24’SIGIR [130] CLIP(ViT-H/14) $51.90$ $74.37$ $53.57$ $73.21$ $58.48$ $79.23$ $54.65$ $75.60$
FashionViL24’ECCV [41] ResNet50 BERT $33.47$ $59.94$ $25.17$ $50.39$ $34.98$ $60.79$ $31.21$ $57.04$
BLIP4CIR224’TMLR [215] BLIP(ViT-B) BLIP $43.78$ $71.34$ $50.15$ $71.25$ $55.23$ $76.80$ $51.17$ $73.13$
CompoDiff23’TMLR [98] CLIP(ViT-L/14) $35.53$ $49.56$ $40.88$ $53.06$ $41.15$ $54.12$ $39.05$ $52.34$
CompoDiff23’TMLR [98] CLIP(ViT-G/14) $38.39$ $51.03$ $41.68$ $56.02$ $45.70$ $57.32$ $39.81$ $51.90$
SDQUR24’TCSVT [65] BLIP2 $49.93$ $73.33$ $56.87$ $76.50$ $59.66$ $79.25$ $55.49$ $76.36$
AlRet24’TMM [39] CLIP LSTM $35.75$ $60.56$ $37.02$ $60.05$ $42.25$ $67.52$ $38.20$ $62.82$
NSFSE24’TMM [47] ResNet50 BiGRU $29.62$ $54.41$ $22.96$ $45.93$ $31.08$ $57.01$ $27.84$ $52.39$
NSFSE24’TMM [47] ResNet152 BiGRU $31.12$ $55.73$ $24.58$ $45.85$ $31.93$ $58.37$ $29.17$ $53.24$
MANME24’TCSVT [53] ResNet50 BiGRU $31.26$ $57.66$ $26.37$ $47.94$ $36.33$ $59.31$ $29.95$ $54.90$
@@
PALAVAR22’ECCV [239] CLIP(ViT-B/32) $17.25$ $35.94$ $21.49$ $37.05$ $20.55$ $38.76$ $19.76$ $37.25$
Pic2Word23’CVPR [12] CLIP(ViT-L/14) $20.00$ $40.20$ $26.20$ $43.60$ $27.90$ $47.40$ $24.70$ $43.70$
SEARLE23’ICCV [23] CLIP(ViT-B/16) $18.54$ $38.51$ $24.44$ $41.61$ $25.70$ $46.46$ $22.89$ $42.53$
SEARLE-XL23’ICCV [23] CLIP(ViT-L/14) $20.48$ $43.13$ $26.86$ $45.58$ $29.32$ $49.97$ $25.56$ $46.23$
MTCIR23’Arxiv [89] CLIP(ViT-L/14) $28.11$ $51.12$ $38.63$ $58.51$ $39.42$ $62.68$ $35.39$ $57.44$
KEDs24’CVPR [81] CLIP(ViT-L/14) $21.7$ $43.8$ $28.9$ $48.0$ $29.9$ $51.9$ $26.8$ $47.9$
PM24’Arxiv [240] CLIP(ViT-L/14) $27.1$ $43.8$ $21.4$ $41.7$ $28.9$ $47.3$ $25.8$ $44.2$
FTI4CIR24’SIGIR [79] CLIP(ViT-L/14) $24.39$ $47.84$ $31.35$ $50.59$ $32.43$ $54.21$ $29.39$ $50.88$
CIReVL24’ICLR [74] CLIP(ViT-L/14) $24.79$ $44.76$ $29.49$ $47.40$ $31.36$ $53.65$ $28.55$ $48.57$
CIReVL24’ICLR [74] CLIP(ViT-G/14) $27.07$ $49.53$ $33.71$ $51.42$ $35.80$ $56.14$ $32.19$ $52.36$
LinCIR24’CVPR [82] CLIP(ViT-L/14) $20.92$ $42.44$ $29.10$ $46.81$ $28.81$ $50.18$ $26.28$ $46.49$
LinCIR24’CVPR [82] CLIP(ViT-G/14) $38.08$ $60.88$ $46.76$ $65.11$ $50.48$ $71.09$ $45.11$ $65.69$
iSEARLE-XL24’Arxiv [77] CLIP(ViT-L/14) $22.51$ $46.36$ $28.75$ $47.84$ $31.31$ $52.68$ $27.52$ $48.96$
RTD(SEARLE)24’Arxiv [83] CLIP(ViT-B/32) $20.72$ $43.13$ $26.69$ $44.31$ $26.67$ $48.75$ $24.7$ $45.4$
RTD((LinCIR)24’Arxiv [83] CLIP(ViT-L/14) $24.49$ $48.24$ $32.83$ $50.44$ $33.4$ $54.56$ $30.24$ $51.08$
ISA(Sym)24’ICLR [80] BLIP $24.69$ $43.88$ $30.79$ $50.05$ $33.91$ $53.65$ $29.79$ $49.19$
ISA(Asy)24’ICLR [80] EfficientNet+BLIP $25.33$ $46.26$ $30.03$ $48.58$ $33.45$ $53.8$ $29.6$ $49.54$
LDRE24’SIGIR [76] CLIP(ViT-L/14) $22.93$ $46.76$ $31.04$ $51.22$ $31.57$ $53.64$ $28.51$ $50.54$
LDRE24’SIGIR [76] CLIP(ViT-G/14) $26.11$ $51.12$ $35.94$ $58.58$ $35.42$ $56.67$ $32.49$ $55.46$
Context-I2W24’AAAI [78] CLIP(ViT-L/14) $23.1$ $45.3$ $29.7$ $48.6$ $30.6$ $52.9$ $27.8$ $48.9$
Slerp+TAT24’ECCV [88] CLIP(ViT-L/14) $23.35$ $45.12$ $29.94$ $46.47$ $31.97$ $51.20$ $28.32$ $47.00$
Slerp+TAT24’ECCV [88] BLIP(ViT-L/16) $29.15$ $50.62$ $32.14$ $51.62$ $37.02$ $57.73$ $32.77$ $53.32$
Denoise-I2W24’Arxiv [141] CLIP(ViT-L/14) $24.4$ $47.8$ $30.9$ $49.8$ $31.6$ $54.1$ $29.0$ $50.6$
MoTaDual(LinCIR)24’Arxiv [86] CLIP-L $28.94$ $49.43$
InstructCIR124’Arxiv [241] CLIP(ViT-L/14) $28.15$ $49.38$ $32.24$ $52.11$ $37.26$ $56.13$ $32.55$ $52.54$
DeG25’Arxiv [140] CLIP(ViT-L/14) $24.4$ $46.5$ $30.7$ $50.3$ $31.6$ $52.0$ $28.9$ $49.6$
Slerp+24’Arxiv [90] BLIP $31.78$ $54.05$ $37.73$ $56.82$ $41.36$ $62.37$ $36.96$ $57.74$
InstructCIR225’Arxiv [143] CLIP(ViT-L/14) $37.32$ $56.84$
TSCIR25’Arxiv [139] CLIP(ViT-L/14) $24.14$ $46.80$ $31.01$ $50.05$ $32.94$ $54.26$ $29.37$ $50.37$
SEIZE24’ACMMM [137] CLIP(ViT-L/14) $30.93$ $50.76$ $33.04$ $53.22$ $35.57$ $58.64$ $33.18$ $54.21$
SEIZE24’ACMMM [137] CLIP(ViT-G/14) $39.61$ $61.02$ $43.60$ $65.42$ $45.94$ $71.12$ $43.05$ $65.85$
PrediCIR25’CVPR [138] CLIP(ViT-L/14) $25.4$ $49.5$ $31.8$ $52.0$ $33.1$ $55.4$ $30.1$ $52.3$
PrediCIR25’CVPR [138] CLIP(ViT-G/14) $39.7$ $62.4$ $48.2$ $67.4$ $53.7$ $73.6$ $47.2$ $67.8$
CoLLM25’CVPR [142] CLIP(ViT-L/14) $30.1$ $49.5$
@@
CompoDiff(ST18M)23’TMLR [98] CLIP(ViT-L/14) $32.24$ $46.27$ $37.69$ $49.08$ $38.12$ $50.57$ $36.02$ $48.64$
PTG(+SPRC)24’Arxiv [92] $31.10$ $51.90$
CASE24’AAAI [96] BLIP $47.44$ $69.36$ $48.48$ $70.23$ $50.18$ $72.24$ $48.79$ $70.68$
HyCIR(CC3M+syn)24’Arxiv [94] CLIP $19.98$ $40.8$ $27.62$ $44.94$ $28.14$ $47.67$ $25.25$ $44.47$
HyCIR(CC3M+syn)24’Arxiv [94] BLIP $18.88$ $34.5$ $22.52$ $37.58$ $22.13$ $40.33$ $21.18$ $37.47$
VDG(+DF’se)24’CVPR [95] BLIP $47.10$ $69.10$ $49.95$ $69.96$ $53.90$ $74.35$ $50.32$ $71.14$
VDG(+FIQ’se)24’CVPR [95] BLIP $47.89$ $69.81$ $51.36$ $71.08$ $53.29$ $74.65$ $50.85$ $71.85$
SCOT25’Arxiv [144] BLIP(ViT-L/16) $26.42$ $49.23$ $30.91$ $49.65$ $34.72$ $55.12$ $30.68$ $51.33$
SCOT25’Arxiv [144] BLIP2(ViT-G/14) $32.78$ $55.91$ $41.42$ $61.09$ $41.15$ $63.10$ $38.45$ $60.03$
InstructCIR225’Arxiv [143] CLIP(ViT-L/14) $49.03$ $70.96$
TSCIR25’Arxiv [139] CLIP(ViT-L/14) $27.22$ $50.87$ $33.71$ $53.43$ $34.73$ $57.22$ $31.88$ $53.84$
MRA-CIR25’Arxiv [150] BLIP2(ViT-L/14) $31.87$ $54.23$ $40.43$ $60.20$ $41.25$ $62.51$ $37.85$ $58.98$
IP-CIR(LDRE)25’CVPR [149] CLIP(ViT-G/14) $39.02$ $61.03$ $48.04$ $66.68$ $50.18$ $71.14$ $45.74$ $66.28$
CIG-XL(SEARLE)25’CVPR [148] CLIP(ViT-B/32) $17.74$ $39.86$ $24.58$ $41.41$ $25.65$ $46.35$ $22.99$ $42.54$
CIG-XL(LinCIR)25’CVPR [148] CLIP(ViT-L/14) $21.27$ $43.98$ $28.66$ $47.20$ $29.83$ $50.28$ $26.59$ $47.15$
CoLLM25’CVPR [142] CLIP(ViT-L/14) $32.9$ $54.2$
CoLLM25’CVPR [142] BLIP(ViT-L/16) $39.1$ $60.7$